OpenStudy (anonymous):

1. $210.00 increased by 40% is what percent of $210.00? 2. How much money is a $210.00 increased by 40%?

OpenStudy (anonymous):

i need help with bot of the questions

OpenStudy (texaschic101):

210 increased by 40% = 210 + 0.40(210)......can you do the math for this one ?

OpenStudy (ahvei):

do you know the answer now

OpenStudy (anonymous):

i tried but i dont think the answer is correct

OpenStudy (ahvei):

whats your answer

OpenStudy (anonymous):

i got 45%

OpenStudy (anonymous):

ik that the 2 one is 224

OpenStudy (texaschic101):

210 + 0.40(210) = 210 + 84 = 294....so $210 increased by 40% is $ 294 now we need to find out : 294 is what percent of 210 all percent problems can be solved by using the percent formula: is/of = %/100 now we need to find out : 294 is what % of 210 is = 294 of = 210 % = x (I just gave it a variable because this is what we need to find now we sub 294/210 = x/100 --- this is a proportion, so we cross multiply 210x = 29400 x = 29400/210 x = 140 so....answer for 1st one is 140% answer for 2nd one is $ 294